OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (lgbasallote):

use distributive property to simplify the equation. what do you get?

OpenStudy (anonymous):

when i add the like terms do i then put the equation like: \[-6x+-9+8\]?

OpenStudy (lgbasallote):

yes. however it's not an "equation" since there's no equal sign

OpenStudy (anonymous):

-3(x+3)=-3(x+1)-5 multiply -3 to everything inside the parenthesis on the left side -3x-9=-3(x+1)-5 follow the same procedure on the right side -3x-9=-3x-3-5 pair like terms (remember when moving on the opposite sides the sign changes) -3x+3x=9-3-5 answer? 0=1 (theres no such thing and therefore can not be solved)

OpenStudy (lgbasallote):

hmm...i think i overlooked that

OpenStudy (anonymous):

so there IS no solution?

OpenStudy (anonymous):

Yes there is no solution

OpenStudy (anonymous):

so thats your answer